# ROTI-GCV: Generalized Cross-Validation for right-ROTationally Invariant Data

### v2

@@ 5.2 Experiments under Signal-PC alignment
(d) Residualized returns
Figure 2: As in Figure 1, curves show the risk and estimated risk as a function of $\lambda$. (2(a)): $\mathbf{X}$ has autocorrelated rows, with $\rho=0.8$. The top 10 eigenvectors ($\mathcal{J}_{c}=[10]$) are coupled, so that the top 10 eigenvectors of the test set are equal to those of $\mathbf{X}$. $\boldsymbol{\beta}=\boldsymbol{\beta}^{\prime}+\sqrt{n}\sum_{i=1}^{10}\frac{i}{10}\mathbf{o}_{i}$, with $\|\boldsymbol{\beta}^{\prime}\|^{2}=\sqrt{n}$, $\sigma^{2}=1$, $n=p=1000$. (2(b)) i.i.d. rows drawn from a Gaussian mixture, i.e. $\mathbf{x}_{i}$ is drawn from $\frac{1}{2}\mathsf{N}(\mathbf{3},\mathbf{I}_{p})+\frac{1}{2}\mathsf{N}(-\mathbf{3},\mathbf{I}_{p})$. When computing ROTI-GCV, we set $\mathcal{J}_{c}=\{1\}$. (2(c)): i.i.d. rows drawn from an equicorrelated Gaussian, i.e. $\mathbf{x}_{i}\sim\mathsf{N}(0,\boldsymbol{\Sigma})$, where $\Sigma_{ij}=\mathds{1}(i=j)+\rho\mathds{1}(i\neq j)$. We set $\rho=0.5$. When computing ROTI-GCV, we set $\mathcal{J}_{c}=[10]$. (2(d)): 30 minute residualized returns sampled every 1 minute;
We now consider settings where the signal $\boldsymbol{\beta}$ aligns with the eigenvectors $\mathbf{O}$. The relevant method here is aROTI-GCV (Algorithm 2). In Figure 2(a), we show the performance of aROTI-GCV when we set the top eigenvectors of the test set equal to those of the training set, matching the coupled eigenvectors condition (Assumption A9). We further construct the signal such that it aligns with the top 10 eigenvectors. As the conditions match our theory, we see aROTI-GCV performs well. The setting of Figure 2(a) might look contrived, but we next provide examples of well-known distributions where such structures naturally arise (Figures 2(b) and 2(c)).
[…]
@@ 5.3 Experiments on semi-real data
In our semi-real experiments, we use real data for the designs, but generate the signals and responses ourselves. Knowing the true signal enables us to calculate the population risk of any estimator, allowing us to evaluate the performance of our GCV metrics accurately. The designs in our experiment consist of residualized returns from 493 stocks retrieved from Polygon API [^2] We remove the top 8 right singular vectors from data matrix $\mathbf{X}$. This is analogous to residualization of returns, a common practice in financial modeling. We compute the factors to delete by taking the top PCs of the covariance matrix computed using both test and train data; PCA is crude/weak as a factor model and we have only 8 factors; hence it is more comparable to compute PCs looking at both train and test data. See appendix B.2 for more details on residualization.. There is one row for each minute, and each row contains the residualized return for that stock over the last 30 minutes. As a result of sharing 29 minutes of returns, consecutive rows are heavily correlated. The regression corresponds to the known practice of replicating the residualized returns of an unknown asset $\mathbf{y}$ using a portfolio of stocks $\hat{\boldsymbol{\beta}}$ (with returns $\mathbf{X}\hat{\boldsymbol{\beta}}$). The loss curves produced by ROTI-GCV better capture the true loss curve, and that the estimated risk values (ER) provide superior estimates of the true out-of-sample risk compared to LOOCV and GCV.
@@ 6 Discussion and Conclusion
Takeaways: We give a precise analysis of ridge regression for right-rotationally invariant designs. These designs depart crucially from the existing framework of i.i.d. samples from distributions with light enough tails, instead allowing for structured dependence and heavy tails. We then characterize GCV in this setting, finding that it is inconsistent, and introduce ROTI-GCV, a consistent alternative; various synthetic and semi-real experiments illustrate the accuracy of our method.
As a byproduct of our analysis, we produce new estimators for signal-to-noise ratio ($r^{2}$) and noise variance ($\sigma^{2}$) robust to these deviations from the i.i.d. assumptions. At the core, our method relies on noticing that the relation given by Lemma 1 yields a different estimating equation for every choice of $\lambda$; using this to generate a series of estimating equations, we obtain stable estimators of $r^{2},\sigma^{2}$. This core idea applies whenever an analogue of Lemma 1 can be derived, giving it reach beyond ridge regression. As future steps, extending our method to other settings such as generalized linear models, kernel ridge regression, regularized random features regression would be important directions. Furthermore, recent advances in [8, 41] show that right-rotationally invariant distributions have a broad universality class – in fact as long as the singular vectors of a design matrix satisfy certain generic structure, properties of learning algorithms under such settings are well captured by results derived under the right-rotationally invariant assumption. In particular, the universality of the in-sample risk of ridge regression is already established in these works, and we believe such universality also holds for the out-of-sample risk; thus, our analysis should extend to the entire universality class of these designs, which is remarkably broad. This promises that our approach can have wide reach, allowing reliable hyperparameter tuning for a broad class of covariate distributions.
@@
Appendix A Deferred Proofs
A.1 Proof of Lemma 6
@@
=$\stackrel{{\scriptstyle(*)}}{{=}}\frac{1}{p}\mathbf{I}_{p}\sum_{i=1}^{n\wedge p}\mathbb{E}[D_{ii}^{2}]=\mathsf{Tr}(\mathbb{E}[\mathbf{D}^{\top}\mathbf{D}])/p\cdot\mathbf{I}_{p}$
where ()(*)( ∗ ) follows from $\mathbf{I}_{p}=\mathbb{E}[\mathbf{O}^{\top}\mathbf{O}]=\sum_{i=1}^{p}\mathbb{E}[\mathbf{o}_{i}\mathbf{o}_{i}^{\top}]$ and the exchangeability of the $\mathbf{o}_{i}$ (which is derived from the fact that $\mathbb{O}(p)$ contains the permutation matrices). ∎
A.2 Useful Results
@@
Isoperimetric inequalities imply that vectors uniformly distributed on $S^{n-1}$ have sub-Gaussian tails around their median (see [32]). Explicitly, it is stated that for $\mathbf{v}\sim\mathrm{Unif}(S^{n-1})$, one has that for any $1$-Lispchitz function $f:\mathbb{R}^{n}\to\mathbb{R}$, one has
@@
where $M$ denotes the median of $f(X)$ (i.e. $\min(\mathbb{P}(f(\mathbf{v})\geq M),\mathbb{P}(f(\mathbf{v})\leq M))\geq\frac{1}{2}$). We now use [5, Theorem 1.2], which shows that having sub-Gaussian tails with respect to the median is equivalent to having sub-Gaussian tails with respect to the mean, with different constants. Thus, it holds that for any $1$-Lipschitz function $f$, one has
@@
Note that this bound is vacuous whenever $14e^{-nt^{2}/32}>$ 1, and hence we can rewrite this as
@@
Finally, [1, Theorem 2.3] implies that random variables satisfying such types of concentration inequalities must in turn satisfy a Hanson-Wright type inequality. That is, Eq. (16) implies that, for fixed matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$, one has
@@
The first equality follows from the cyclic property of trace. The second can be seen to follow from the observation used in the proof of Lemma 6 that $\mathbb{E}[\mathbf{o}_{i}\mathbf{o}_{i}^{\top}]=\frac{1}{p}\mathbf{I}_{p}$ and the fact that columns of a Haar distributed matrix are uniform on the sphere (see e.g. [28]). Alternatively, recall that for $\mathbf{g}\sim\mathsf{N}(0,\mathbf{I}_{n})$, $\frac{\mathbf{g}}{\|\mathbf{g}\|}\perp\!\!\!\perp\|\mathbf{g}\|$ from [27, Theorem 2.7.1]. Hence
@@
Let $\mathbf{v}_{n}$ be a sequence of random vectors, where $\mathbf{v}_{n}$ is uniformly distributed on $S^{n-1}$. Let $\mathbf{u}_{n}$ be another sequence of random vectors independent of $\mathbf{v}_{n}$, where lim $\limsup\|\mathbf{u}_{n}\|<C$ for some constant $C$ almost surely. Then $\mathbf{v}_{n}^{\top}\mathbf{u}_{n}\xrightarrow{a.s.}0$.

### v3

@@ 5.2 Experiments under Signal-PC alignment
(d) Speech Data without Signal-PC alignment
(e) Speech Data with Signal-PC alignment
(f) Residualized returns
Figure 2: As in Figure 1, curves show the risk and estimated risk as a function of $\lambda$. (2(a)): $\mathbf{X}$ has autocorrelated rows, with $\rho=0.8$. The top 10 eigenvectors ($\mathcal{J}_{c}=[10]$) are coupled, so that the top 10 eigenvectors of the test set are equal to those of $\mathbf{X}$. $\boldsymbol{\beta}=\boldsymbol{\beta}^{\prime}+\sqrt{n}\sum_{i=1}^{10}\frac{i}{10}\mathbf{o}_{i}$, with $\|\boldsymbol{\beta}^{\prime}\|^{2}=\sqrt{n}$, $\sigma^{2}=1$, $n=p=1000$. (2(b)) i.i.d. rows drawn from a Gaussian mixture, i.e. $\mathbf{x}_{i}$ is drawn from $\frac{1}{2}\mathsf{N}(\mathbf{3},\mathbf{I}_{p})+\frac{1}{2}\mathsf{N}(-\mathbf{3},\mathbf{I}_{p})$. When computing ROTI-GCV, we set $\mathcal{J}_{c}=\{1\}$. (2(c)): i.i.d. rows drawn from an equicorrelated Gaussian, i.e. $\mathbf{x}_{i}\sim\mathsf{N}(0,\boldsymbol{\Sigma})$, where $\Sigma_{ij}=\mathds{1}(i=j)+\rho\mathds{1}(i\neq j)$. We set $\rho=0.5$. When computing ROTI-GCV, we set $\mathcal{J}_{c}=[10]$. (2(d)): speech data with $\boldsymbol{\beta}$ sampled uniformly from sphere, with $r^{2}=\sigma^{2}=1$, $n=p=400$; we choose $\mathcal{J}_{c}=[3]$. (2(e)): speech data with $\boldsymbol{\beta}=\boldsymbol{\beta}^{\prime}+\tfrac{\sqrt{n}}{2}\sum_{i=1}^{5}\mathbf{o}_{i}$; again $r^{2}=\sigma^{2}=1$ and $n=p=400$. We choose $\mathcal{J}_{c}=[3]$ and $\mathcal{J}_{a}=[5]$ (discussion in Section 5.3.1). (2(f)): 30 minute residualized returns sampled every 1 minute; $\boldsymbol{\beta}$ sampled uniformly from sphere.
We now consider settings where the signal $\boldsymbol{\beta}$ aligns with the eigenvectors $\mathbf{O}$. The relevant method here is aROTI-GCV (Algorithm 2). In Figure 2(a), we show the performance of aROTI-GCV when we set the top eigenvectors of the test set equal to those of the training set, matching the coupled eigenvectors condition (Assumption A9). We further construct the signal such that it aligns with the top 10 eigenvectors. As the conditions match our theory, we see aROTI-GCV performs well. The setting of Figure 2(a) might look contrived, but we next provide examples of well-known distributions where such structures naturally arise (Figures 2(b) and 2(c)). In Figure 2(b), the data is drawn from a mixture of two Gaussians, one centered at $\mathbf{3}=[3,3,\dots,3]$ and the other centered at $\mathbf{3}$, each with identity covariance. In this setting, the top eigenvector is very close to $\mathbf{3}$ and emerges from the bulk. Furthermore, projected into the subspace $\mathrm{span}(\mathbf{3})^{\perp}$, the data is a standard Gaussian and thus right-rotationally invariant. As seen, our method works under this type of covariance structure. Shown in Figure 2(c), we find similar results when each row of $\mathbf{X}$ is drawn from an equicorrelated Gaussian, i.e., $\mathbf{x}_{i}\sim\mathsf{N}(0,\boldsymbol{\Sigma})$ where $\Sigma_{ij}=\rho^{1-\mathds{1}(i=j)}$. Both examples mirror the setting of 2(a) in that they are not right-rotationally invariant, as the test and train tests vary in one direction more than the others. However, we see that the modifications in aROTI-GCV enable the method to still succeed. Note that in both of these examples, the top eigenvalue diverges at the rate $O(n)$, which violates our Assumption A11 as well as the bounded spectrum condition (Patil et al., , 2021, Assumption 3), meaning neither cross-validation method is provably valid in this setting. However, we observe that ROTI-GCV still performs reasonably well.
@@ 5.3 Experiments on semi-real data
In our semi-real experiments, we use real data for the designs, but generate the signals and responses ourselves. Knowing the true signal enables us to calculate the population risk of any estimator, allowing us to evaluate the performance of our GCV metrics accurately. For such experiments, we must verify whether the assumptions of our theory are met. To this end, we provide a series of diagnostics to check whether the most important assumptions hold, including to select the index sets $\mathcal{J}_{a}$ and $\mathcal{J}_{c}$. We briefly list these here, and defer more details to Appendix 9, including examples of running these diagnostics for multiple datasets.
(i) regularity of singular value distributions (Assumptions A2, A11). Plot histogram of singular values to check for outliers; see Figure 3.
(ii) signal-PC alignment: relationship between $\boldsymbol{\beta}$ and $\{\mathbf{o}_{i}\}_{i=1}^{p}$ (Assumptions A6, A8). Li and Sur, (2023) provide a hypothesis testing framework for identifying $\mathcal{J}_{a}$, wherein one computes a $p$-value per eigenvector to test whether or not it is aligned; the Benjamini-Hochberg procedure can then be used to control the false discovery rate. See Table 2.
(iii) coupling: relationship between $\{\mathbf{o}_{i}\}_{i=1}^{p}$ and $\{\widetilde{\mathbf{o}}_{j}\}_{i=1}^{p}$ (Assumptions A7, A9). We compute overlaps $\sqrt{p}\langle\mathbf{o}_{i},\widetilde{\mathbf{o}}_{j}\rangle$ for the top singular vectors. Note if $\mathbf{o}_{i}~\perp\!\!\!\perp~\widetilde{\mathbf{o}}_{j}$ (as under A7), then $\sqrt{p}\langle\mathbf{o}_{i},\widetilde{\mathbf{o}}_{j}\rangle\to\mathsf{N}(0,1)$. If $\sqrt{p}\langle\mathbf{o}_{i},\widetilde{\mathbf{o}}_{j}\rangle$ is large relative to this null distribution, $\mathbf{o}_{i}$, $\widetilde{\mathbf{o}}_{j}$ should be coupled. See Table 1. We leave development of a formal hypothesis test to future work.
5.3.1 Speech data
The designs we use consist of speech data sourced from OpenML (Le, , 2017). Each row has dimension 400 and consists of an i-vector (a type of featurization for audio data, see e.g. Ibrahim and Ramli, (2018)) of a speech segment. In this first example, we fully illustrate all the proposed diagonstics; as such, we consider both when $\boldsymbol{\beta}$ is uniformly drawn on the sphere, so there is no signal-PC alignment, and when signal-PC alignment is present, to illustrate how to choose $\mathcal{J}_{a}$.
We first examine the distribution of singular values in Figure 3, finding there are no outlier values. Next, we compute overlaps of test and train eigenvectors, shown in Table 1, finding that the top 3 eigenvectors of each have strong overlaps. This motivates setting $\mathcal{J}_{c}=[3]$ [^4] One can further couple the next few eigenvectors (e.g. $\mathbf{o}_{6}$ with $\widetilde{\mathbf{o}}_{4}$ and so on), but we observe this has no impact on the results.. The results for this setting when $\boldsymbol{\beta}$ is drawn uniformly from the sphere are shown in Figure 2(d), where we see successfully capture the true loss curve and that the estimated risk values are superior to those given by GCV and LOOCV.
Next, in the setting with added signal-PC alignment, we instead take $\boldsymbol{\beta}=\boldsymbol{\beta}^{\prime}+\tfrac{\sqrt{n}}{2}\sum_{i=1}^{5}\mathbf{o}_{i}$, meaning the aligned set is $\mathcal{J}_{a}=[5]$. We use the hypothesis testing framework of Li and Sur, (2023) and compute Benjamini-Hochberg-adjusted $p$-values for alignment, shown in Table 2; here, we clearly identify the set of aligned eigenvectors. The resulting loss curves for this setting are then shown in Figure 2(e).
Figure 3: Singular value distribution for speech data.
Table 1: Overlaps for speech data setting; Cell $(i,j)$ contains $\sqrt{p}|\langle\mathbf{o}_{i},\widetilde{\mathbf{o}}_{j}\rangle|$.
$\widetilde{\mathbf{o}}_{1}$ $\widetilde{\mathbf{o}}_{2}$ $\widetilde{\mathbf{o}}_{3}$ $\widetilde{\mathbf{o}}_{4}$ $\widetilde{\mathbf{o}}_{5}$ $\widetilde{\mathbf{o}}_{6}$ $\widetilde{\mathbf{o}}_{7}$ $\widetilde{\mathbf{o}}_{8}$ $\widetilde{\mathbf{o}}_{9}$ $\widetilde{\mathbf{o}}_{10}$
$\mathbf{o}_{1}$ 17.30 2.72 0.63 2.01 1.55 1.02 0.30 0.96 1.70 0.27
$\mathbf{o}_{2}$ 0.83 15.24 0.68 1.63 1.34 1.22 0.89 1.42 0.27 2.19
$\mathbf{o}_{3}$ 1.08 1.93 10.93 1.60 0.64 2.02 4.63 0.35 4.14 2.75
$\mathbf{o}_{4}$ 0.29 2.78 2.92 0.22 3.98 6.63 2.98 1.65 0.54 1.56
$\mathbf{o}_{5}$ 1.27 2.70 4.02 2.14 5.11 1.29 4.34 4.72 2.95 2.24
$\mathbf{o}_{6}$ 1.19 0.56 4.82 5.95 0.37 3.83 1.37 0.87 4.98 1.04
$\mathbf{o}_{7}$ 2.21 0.70 2.14 2.65 4.84 1.77 0.54 1.00 1.65 2.53
$\mathbf{o}_{8}$ 0.46 0.15 2.96 0.16 1.66 4.49 0.22 0.40 1.82 0.10
$\mathbf{o}_{9}$ 0.30 0.40 0.72 0.01 1.63 1.39 3.86 4.53 1.12 2.57
$\mathbf{o}_{10}$ 1.14 2.51 0.90 0.77 0.37 0.79 2.44 1.83 1.81 2.34
Table 2: Benjamini-Hochberg adjusted $p$-values for alignment.
$\mathbf{o}_{1}$ $\mathbf{o}_{2}$ $\mathbf{o}_{3}$ $\mathbf{o}_{4}$ $\mathbf{o}_{5}$
pp 0.000 0.000 0.000 0.000 0.000
$\mathbf{o}_{6}$ $\mathbf{o}_{7}$ $\mathbf{o}_{8}$ $\mathbf{o}_{9}$ $\mathbf{o}_{10}$
pp 0.934 0.588 0.651 0.913 0.395
5.3.2 Residualized returns
@@
We next consider designs consisting of residualized returns from 493 stocks retrieved from Polygon API [^5] We remove the top 8 right singular vectors from data matrix $\mathbf{X}$. This is analogous to residualization of returns, a common practice in financial modeling. We compute the factors to delete by taking the top PCs of the covariance matrix computed using both test and train data; PCA is crude/weak as a factor model and we have only 8 factors; hence it is more comparable to compute PCs looking at both train and test data. See Appendix 8.2 for more details on residualization. (Polygon, , 2024). There is one row per minute, and each contains the residualized return for each stock over the last 30 minutes. Since they share 29 minutes of returns, consecutive rows are heavily correlated. The regression corresponds to the known practice of replicating the residualized returns of an unknown asset $\mathbf{y}$ using a portfolio of stocks $\hat{\boldsymbol{\beta}}$ (with returns $\mathbf{X}\hat{\boldsymbol{\beta}}$). The diagnostic plots in this case are tame, and we require no coupling; these are deferred to Appendix 9. The loss curves produced by ROTI-GCV better capture the true loss curve, and that the estimated risk values (ER) provide superior estimates of the true out-of-sample risk compared to LOOCV and GCV.
@@ 6 DISCUSSION AND CONCLUSION
Takeaways: We give a precise analysis of ridge regression for right-rotationally invariant designs. These designs depart crucially from the existing framework of i.i.d. samples from distributions with light enough tails, instead allowing for structured dependence and heavy tails. We then characterize GCV in this setting, finding that it is inconsistent, and introduce ROTI-GCV, a consistent alternative; various synthetic and semi-real experiments illustrate the accuracy of our method. As a byproduct of our analysis, we produce new estimators for signal-to-noise ratio ($r^{2}$) and noise variance ($\sigma^{2}$) robust to these deviations from the i.i.d. assumptions. At the core, our method relies on noticing that the relation given by Lemma 1 yields a different estimating equation for every choice of $\lambda$; using this to generate a series of estimating equations, we obtain stable estimators of $r^{2},\sigma^{2}$. This core idea applies whenever an analogue of Lemma 1 can be derived, giving it reach beyond ridge regression. As future steps, extending our method to other settings such as generalized linear models, kernel ridge regression, regularized random features regression would be important directions. Furthermore, recent advances in Dudeja et al., (2023); Wang et al., (2023) show that right-rotationally invariant distributions have a broad universality class – in fact as long as the singular vectors of a design matrix satisfy certain generic structure, properties of learning algorithms under such settings are well captured by results derived under the right-rotationally invariant assumption. In particular, the universality of the in-sample risk of ridge regression is already established in these works, and we believe such universality also holds for the out-of-sample risk; thus, our analysis should extend to the entire universality class of these designs, which is remarkably broad. This promises that our approach can have wide reach, allowing reliable hyperparameter tuning for a broad class of covariate distributions.
@@
Appendix
7 Deferred Proofs
7.1 Proof of Lemma 6
@@
$\stackrel{{\scriptstyle(*)}}{{=}}\frac{1}{p}\mathbf{I}_{p}\sum_{i=1}^{n\wedge p}\mathbb{E}[D_{ii}^{2}]=\mathsf{Tr}(\mathbb{E}[\mathbf{D}^{\top}\mathbf{D}])/p\cdot\mathbf{I}_{p}$
where ()(*) follows from $\mathbf{I}_{p}=\mathbb{E}[\mathbf{O}^{\top}\mathbf{O}]=\sum_{i=1}^{p}\mathbb{E}[\mathbf{o}_{i}\mathbf{o}_{i}^{\top}]$ and the exchangeability of the $\mathbf{o}_{i}$ (which is derived from the fact that $\mathbb{O}(p)$ contains the permutation matrices). ∎
7.2 Useful Results
@@
Isoperimetric inequalities imply that vectors uniformly distributed on $S^{n-1}$ have sub-Gaussian tails around their median (see Schechtman, (2014)). Explicitly, it is stated that for $\mathbf{v}\sim\mathrm{Unif}(S^{n-1})$, one has that for any $1$-Lispchitz function $f:\mathbb{R}^{n}\to\mathbb{R}$, one has
@@
where $M$ denotes the median of $f(X)$ (i.e. $\min(\mathbb{P}(f(\mathbf{v})\geq M),\mathbb{P}(f(\mathbf{v})\leq M))\geq\frac{1}{2}$). We now use (Chazottes, , 2022, Theorem 1.2), which shows that having sub-Gaussian tails with respect to the median is equivalent to having sub-Gaussian tails with respect to the mean, with different constants. Thus, it holds that for any $1$-Lipschitz function $f$, one has
@@
Note that this bound is vacuous whenever $4e^{-nt^{2}/32}>1$, and hence we can rewrite this as
@@
Finally, (Adamczak, , 2015, Theorem 2.3) implies that random variables satisfying such types of concentration inequalities must in turn satisfy a Hanson-Wright type inequality. That is, Eq. (16) implies that, for fixed matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$, one has
@@
The first equality follows from the cyclic property of trace. The second can be seen to follow from the observation used in the proof of Lemma 6 that $\mathbb{E}[\mathbf{o}_{i}\mathbf{o}_{i}^{\top}]=\frac{1}{p}\mathbf{I}_{p}$ and the fact that columns of a Haar distributed matrix are uniform on the sphere (see e.g. Meckes, (2019)). Alternatively, recall that for $\mathbf{g}\sim\mathsf{N}(0,\mathbf{I}_{n})$, $\frac{\mathbf{g}}{\|\mathbf{g}\|}\perp\!\!\!\perp\|\mathbf{g}\|$ from (Mardia et al., , 2024, Theorem 2.7.1). Hence
@@
Let $\mathbf{v}_{n}$ be a sequence of random vectors, where $\mathbf{v}_{n}$ is uniformly distributed on $S^{n-1}$. Let $\mathbf{u}_{n}$ be another sequence of random vectors independent of $\mathbf{v}_{n}$, where $\limsup\|\mathbf{u}_{n}\|<C$ for some constant $C$ almost surely. Then $\mathbf{v}_{n}^{\top}\mathbf{u}_{n}\xrightarrow{a.s.}0$.